OpenStudy (anonymous):

write the following quadratics in the form y=(x-h)^2 +k using completing the squares and hence sketch each function and stating the vertex a) y= x^2 -6x+5 b) y= x^2 +8x -2

OpenStudy (ash2326):

a) y=x^2-6x+5 y=x^2-6x+9-4 y=(x-3)^2+(-4) vertex (3,-4) b) y=x^2+8x+5 y=x^2+8x+16-21 y=(x+4)^2-21 y=(x-(-4))^2-21 vertex (-4,-21)

jhonyy9 (jhonyy9):

b. not is correct

jhonyy9 (jhonyy9):

why -21 ?

jhonyy9 (jhonyy9):

just -11

jhonyy9 (jhonyy9):

right ???

jhonyy9 (jhonyy9):

but in point b. there is +2 and not +5 like last term

OpenStudy (ash2326):

sorry -11

jhonyy9 (jhonyy9):

oh not +5 there is -2

OpenStudy (ash2326):

y=x^2+8x+5 y=x^2+8x+16-11 y=(x+4)^2-11 y=(x-(-4))^2-11 vertex (-4,-11)

OpenStudy (anonymous):

hmmm i think its supposed to be 18?

OpenStudy (anonymous):

woopppss the question b) was supposed to be x^2+8x-2

jhonyy9 (jhonyy9):

you need checking how can you get 8 can us 6+2 or 4+4 or 5+3 or 7+1 or ????

OpenStudy (anonymous):

instead of +5 haha

OpenStudy (ash2326):

sorry kirsten i solved for +5 y=(x-(-4))^2-18 vertex (-4,-18)

jhonyy9 (jhonyy9):

yes and after you check how can you getting the needed number what will result from product of used numbers for 8 and after you will get how many you need subtracting from that

jhonyy9 (jhonyy9):

ok ?

OpenStudy (anonymous):

yapppp :) awesome thank you guys